OpenStudy (anonymous):

how to solve 1/1-cos2x?

OpenStudy (anonymous):

solve what ? ! an equation ?!

OpenStudy (anonymous):

how do you verify that 1/1-cos2x is \[\left(\begin{matrix}1 \\ 2\end{matrix}\right)\] csc\[^{2}\] x

OpenStudy (anonymous):

1/2csc²x

OpenStudy (anonymous):

kk use this \[\cos(2x)=1-2\sin^2(x)\] calculate first 1-2cos(2x) what did you get ?

OpenStudy (anonymous):

1/sin²x?

OpenStudy (anonymous):

sorry ;) 1-cos(2x) = ?! (I' typed 1-2cos(2x))

OpenStudy (anonymous):

i dont get it...

OpenStudy (anonymous):

cos2x= 1-2sin²x

OpenStudy (anonymous):

\[\text{using this } \color{green}{\cos(2x)}=1-2\sin^2(x) \text{ what is } 1-\color{green}{\cos(2x)}?\]

OpenStudy (anonymous):

1-1-sin²x...?

OpenStudy (anonymous):

\[1-(1-2\sin^2(x))\]

OpenStudy (anonymous):

oh okay i got it

OpenStudy (anonymous):

what did you get .? :)

OpenStudy (anonymous):

it would be 1/ 1-(1-2sin²(x))

OpenStudy (anonymous):

But i dont know what to do on the next step.....

OpenStudy (anonymous):

C'moon !

OpenStudy (anonymous):

so what's \[-(1-2\sin^2(x)) \text{ without using this "(" ")"}\]

OpenStudy (anonymous):

-cos2x

OpenStudy (anonymous):

okay i am really lost doesnt it become -2sin²x?

OpenStudy (anonymous):

That's not what I'm asking for ! what's -(a-b) without using this "(" ")"

OpenStudy (anonymous):

-a-b

OpenStudy (anonymous):

no ! -a+b An example -(4-1)=-4+1=-3

OpenStudy (anonymous):

OHHH yes yes yes i got what you mean

OpenStudy (anonymous):

now ! if you are convinced ! what's \[1-(1-2\sin^2(x))\]

OpenStudy (anonymous):

1-1+2sin²x

OpenStudy (anonymous):

which is 2sin²x !!!!???

OpenStudy (anonymous):

yes ! then 1/1-cos(2x) =?

OpenStudy (anonymous):

are you there @lsugano

OpenStudy (anonymous):

um so do we use this identity sinx = 1/cscx

OpenStudy (anonymous):

yes !

OpenStudy (anonymous):

what did you get ?

OpenStudy (anonymous):

i dont know how the answer is 1/2cscx

OpenStudy (anonymous):

is this step correct? http://bamboodock.wacom.com/doodler/2aff60b5-e920-4961-bb39-e9ab34c722b6

OpenStudy (anonymous):

C'mooon ! \[\text{we said that } 1-\cos(2x)=2\sin^2(x)\] then \[\frac{1}{1-\cos(2x)}=\frac{1}{2\sin^2(x)}=\frac{1}{2}(\frac{1}{\sin(x)})^2\] do you see that .

OpenStudy (anonymous):

Soryy :'( I didn't see your response !

OpenStudy (anonymous):

Okayy hahaha i am really sorry i made this very difficult. i am a bad math learner

OpenStudy (anonymous):

BUT i got what this means!!! THank you so much!

OpenStudy (anonymous):

what you did is also right !

OpenStudy (anonymous):

so it becomes 1/2csc²x!!!!

OpenStudy (anonymous):

yes ! \[\frac{1}{2} \csc^2(x) \]

OpenStudy (anonymous):

oohh thank you so much! i couldnt have gone through this without you, seriously thank you thank you!!!

OpenStudy (anonymous):

I checked your profile :D ! WELCOME TO OPENSTUDY ! you're welcome anytime ! (Lisa) :) :)

OpenStudy (anonymous):

Thank you ! yeah, im really new to this so I hope i get used to it!! thank you !!

OpenStudy (anonymous):

Yw :) if you need Anything else just send me the link of your question Or mention me ( @Neemo) !

OpenStudy (anonymous):

Okay thank you very much!

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

kk ! :)

OpenStudy (anonymous):

How do you solve sin⁴a=(sin²a)²

OpenStudy (anonymous):

so far i know that its \[\left(\begin{matrix}1-\cos2a \ \div\ 2\end{matrix}\right)^{2}\]

OpenStudy (anonymous):

kk you mean this \[\sin^4(x)=(\sin^2(x))^2\] ?

OpenStudy (anonymous):

yes!!!

OpenStudy (anonymous):

that's not obvious for you ! \[(a^2)^2=?\]

OpenStudy (anonymous):

(2²)²=16 like that?

OpenStudy (anonymous):

i dont know how variable work with those.....

OpenStudy (anonymous):

maybe ! but ! in general case we have this property \[(a^n)^p=a^{np}\]

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

your example ! \[(2^2)^2=2^{(2).(2)}=2^4=16\] so yes !

OpenStudy (anonymous):

so what do we do next for this one?

OpenStudy (anonymous):

we square 1-cos2x/2 right

OpenStudy (anonymous):

It obvious ! \[(\sin^2(x))^2=(\sin(x))^4=\sin^4(x)\]

OpenStudy (anonymous):

why you want to do that ?!

OpenStudy (anonymous):

the question says write sin⁴x in terms of the first power of one or more cosine functions. using power reducing identites. and sin²x=1-cos2x/2

OpenStudy (anonymous):

ohh ! you didn't say that :) :) ?!

OpenStudy (anonymous):

so since sin⁴x=(sin²x)²=====(1-cos2x/2)²

OpenStudy (anonymous):

im sorry I didn't mention that !!

OpenStudy (anonymous):

No prblem at all ! yes !go on I'm watching you ! :)

OpenStudy (anonymous):

so is it like 1-cos2x² / 4

OpenStudy (anonymous):

remember ! that you 'll go to jail if you ever do that again !this is the property !! \[(a-b)^2=a^2-2ab+b^2 \] and remember \[a^2-b^2=(a-b)(a+b) \text{ \not what you typed } ==////===(a-b)^2\]

OpenStudy (anonymous):

so correct your mistake :) @lsugano

OpenStudy (anonymous):

i will write down on a tablet and could you check if im right?

OpenStudy (anonymous):

I'll be happy :)

OpenStudy (anonymous):

were you able to see it?

OpenStudy (anonymous):

yes ! But It's you shouldn't have \[\cos^2(2x)\]

OpenStudy (anonymous):

sorry for being late :) !

OpenStudy (anonymous):

@lsugano how can you deal with that ?!

OpenStudy (anonymous):

yes! okay i think i got the answer!!

OpenStudy (anonymous):

because you reminded me about (a-b)squared, this equation was really easy to figure out

OpenStudy (anonymous):

can I check it ! caus that isn't supposed to be the final answer ! :)

OpenStudy (anonymous):

3-4cos2x+cos4x/8

OpenStudy (anonymous):

Sorry ! I underestimated you ! you're really smart :) That's correct :) !

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

thank you somuch! you a great teacher!!

OpenStudy (anonymous):

No !usually, I'm a very bad teacher ! just you're the special one ! another question ?

OpenStudy (anonymous):

okay so i got sin²xcos⁴x

OpenStudy (anonymous):

yeah ! what is the question ?!

OpenStudy (anonymous):

its the same thing

OpenStudy (anonymous):

and my first step i did was this

OpenStudy (anonymous):

kk ! go on !

OpenStudy (anonymous):

and i got this one

OpenStudy (anonymous):

correct ! go on !

OpenStudy (anonymous):

are you thre ?! @lsugano

OpenStudy (anonymous):

ssryy im am almost done!!

OpenStudy (anonymous):

could you please check?